# Drop rebound at low Weber number

### v1

@@ 4.3 Maximum spreading and deformation
We next examine how drop deformation depends on the governing dimensionless parameters and relate deformation and spreading. To do this, we define the maximum equatorial deformation $\beta-1$, where $\beta=R_{m}/R$ is the maximum equatorial radius. Similarly, we define maximum vertical deformation $1-\alpha$, where $\alpha=H_{min}/2R$ and $H_{min}$ is the minimum drop height. Lastly, we define the maximum spreading radius as $\beta_{c}=R_{c}/R$.
@@ 4.3.1 Energy argument
with the surface energy of the $l=2$ mode being $E_{2}=\frac{8}{5}\pi\sigma\mathscr{A}_{2}^{\prime 2}$ (Moláček & Bush, 2012). Neglecting gravity ($B\kern-0.50003pto\,^{2}\ll W\kern-1.49994pte\,$) and viscous dissipation ($O\kern-0.50003pth\,\ll 1$), conservation of energy suggests
@@
with the result for $\beta-1$ being identical to that derived by Richard & Quéré (2000) using similar arguments. This derivation also suggests a geometric relationship between $\beta$ and $\alpha$:
@@ 4.3.2 Comparisons
Figure 8: ($a$) Maximum equatorial deformation $\beta-1$ ($\beta=R_{m}/R$) and ($b$) maximum vertical deformation $1-\alpha$ ($\alpha=H_{min}/2R$) plotted against Weber number $W\kern-1.34995pte\,$. Circles represent experiment and stars denote DNS for $B\kern-0.45003pto\,=0.02$. Solid curves show the KM model with $B\kern-0.45003pto\,=0.02$, and the inertio-capillary limit ($O\kern-0.45003pth\,=0$; $B\kern-0.45003pto\,=0$) is represented by black dashed curves. The red dot-dashed curves are predictions from our simplified model (§ 4.3.1). The inset in ($a$) shows our low-viscosity data $O\kern-0.45003pth\,<0.05$ (colored circles) with prior results (diamond (Okumura et al., 2003), pentagon (Clanet et al., 2004), square (Biance et al., 2006), triangle (Laan et al., 2014)). The red dot-dashed curve is equation (30), and the black solid line with slope 1/4 is used to guide the eye. The inset in ($c$) plots $1-\alpha$ against $\beta-1$ for all experiments and the red dot-dashed curve is equation (32). ($c$) Dimensionless spreading area $\beta_{c}^{2}$ against $\beta-1$ for all experiments and DNS results, with equation (33) shown as a red dot-dashed curve. The color bar maps $O\kern-0.45003pth\,$ to color on a logarithmic scale.
@@ 5 Discussion
The coefficient of restitution $\varepsilon$ and contact time $t_{c}$ have different $W\kern-1.49994pte\,$-dependencies at low and high $W\kern-1.49994pte\,$. When $W\kern-1.49994pte\,\gtrapprox 0.5$, the coefficient of restitution $\varepsilon$ decreased as $W\kern-1.49994pte\,$ increased, as observed in prior studies (Biance et al., 2006; Jha et al., 2020). When $W\kern-1.49994pte\,\lessapprox 0.5$, $\varepsilon$ reached a maximum that depends upon $O\kern-0.50003pth\,$ and $B\kern-0.50003pto\,$, and becomes $W\kern-1.49994pte\,$-independent in the inertio-capillary limit. A gravity-induced roll-off in $\varepsilon$ leading to $\varepsilon\rightarrow 0$ is the natural route to rebound suppression when $B\kern-0.50003pto\,$ is finite, with $O\kern-0.50003pth\,$ increasing the critical Weber number $W\kern-1.49994pte\,_{c}$ below which rebound stops. These results differ from the peak and subsequent roll-off in $\varepsilon$ as $W\kern-1.49994pte\,$ decreases reported for $W\kern-1.49994pte\,<1$ rebounds on superhydrophobic substrates (Richard & Quéré, 2000; Wang et al., 2022; Thenarianto et al., 2023). Although these prior experiments were conducted in low $O\kern-0.50003pth\,$ conditions, and in some cases used low $B\kern-0.50003pto\,$ drops, the increased effect of contact line friction at low $W\kern-1.49994pte\,$ was shown to be the primary mechanism for rebound suppression. Consequently, our results show that submillimeter drops, like those from coughing (Dbouk & Drikakis, 2020) or agricultural sprays (Kooij et al., 2018; Makhnenko et al., 2021; Chen & Ashgriz, 2022), may continue bouncing longer than previously thought when the substrate is non-wetting—an important insight into the spread of their contents. The contact time between the drop and substrate is $W\kern-1.49994pte\,$-dependent at low $W\kern-1.49994pte\,$ (Foote, 1975; Okumura et al., 2003), and our results show that it is also $O\kern-0.50003pth\,$-dependent in this regime. Increasing $O\kern-0.50003pth\,$, causes $t_{c}$ to increase at a higher rate as $W\kern-1.49994pte\,$ decreases such that viscosity is most influential on $t_{c}$ in the low-$W\kern-1.49994pte\,$ regime. Since $t_{c}$ is a critical parameter governing drop–substrate interaction in various applications Shiri & Bird (2017); Díaz et al. (2022), this finding has immediate practical implications, and motivates the study of low $W\kern-1.49994pte\,$ rebound using drops with shear-dependent viscosity in future work.
@@
In section 4.2, we connected the unique rebound behavior at low $W\kern-1.49994pte\,$ to the spreading and retraction dynamics. We compared the drop shape during spreading and retraction for three specific $W\kern-1.49994pte\,$ spanning the transition from low $W\kern-1.49994pte\,$ to high $W\kern-1.49994pte\,$, and revealed a nearly symmetric spreading and retraction process for low $O\kern-0.50003pth\,$, low $W\kern-1.49994pte\,$ rebound. As $W\kern-1.49994pte\,$ increased, the spread time decreased relative to the retraction time, increasing asymmetry between spreading and retraction and aligning with the changes in rebound behavior observed as $W\kern-1.49994pte\,$ increases. The mechanism for asymmetry is the limited maximum retraction rate a drop can attain, which is substrate-independent for low $O\kern-0.50003pth\,$ (Bartolo et al., 2005). This maximum retraction rate is not reached during retraction when $W\kern-1.49994pte\,$ is sufficiently small, and a constant proportion of the drops time in contact with the substrate is spent spreading and retracting. In the inertio-capillary limit, the spreading time approximately equals the retraction time, but as $O\kern-0.50003pth\,$ increases the spreading time decreases relative to the retraction time, as shown in figure 7. The symmetry between the spreading and retraction phase in the inertio-capillary limit is lost when $W\kern-1.49994pte\,$ becomes sufficiently large, and our low $O\kern-0.50003pth\,$ data follows $t_{s}/t_{r}\sim W\kern-1.49994pte\,^{-0.26\pm 0.01}$ at high $W\kern-1.49994pte\,$. Considering the time needed for a drop to spread $t_{s}\sim R_{c}/V$, where $R_{c}\sim R\,W\kern-1.49994pte\,^{-1/4}$ from our energy argument (34), and its maximum retraction rate $t_{r}^{-1}\sim t_{\sigma}^{-1}$ (Bartolo et al., 2005), we arrive at $t_{s}/t_{r}\sim W\kern-1.49994pte\,^{-1/4}$, in agreement with our results. Recent studies have noted the spreading and retraction time of a rebounding drop are similar at low $W\kern-1.49994pte\,$ (Lin et al., 2024), and our low-$O\kern-0.50003pth\,$ results support this finding and extend its broader relevance by demonstrating its relation to the overall dynamics of rebound. Additionally, our results reveal that increasing $O\kern-0.50003pth\,$ increases asymmetry, thereby altering the rebound behavior even at low $W\kern-1.49994pte\,$.
@@
Lastly, we examined the drop deformation and spreading in section 4.3, connecting them through energy arguments and demonstrating the accuracy of the KM model for predicting the drop shape during impact. The literature on the maximum horizontal deformation is vast and two competing scaling arguments have gained support: one based on momentum conservation $\beta\sim W\kern-1.49994pte\,^{1/4}$ (Clanet et al., 2004; Bartolo et al., 2005; Biance et al., 2006; Tsai et al., 2011; García-Geijo et al., 2020) and the other arising from energy conservation $\beta\sim We^{1/2}$ (Chandra & Avedisian, 1991; Bennett & Poulikakos, 1993; Okumura et al., 2003; Eggers et al., 2010; Villermaux & Bossa, 2011; Laan et al., 2014; Lee et al., 2016). Despite extensive efforts, the correct scaling remains debated, in part due to most studies using the maximum equatorial radius $\beta=R_{m}/R$ to describe horizontal deformation. However, as discussed in (Josserand & Thoroddsen, 2016), $\beta$ ignores the initial drop radius and is inadequate to discern scaling arguments at low $W\kern-1.49994pte\,$. In this work we used the maximum equatorial deformation $\beta-1$, and energy arguments predict $\beta-1=\left(5/48\right)^{1/2}\,We^{1/2}$, as in Richard & Quéré (2000). This prediction is in excellent agreement with our low $O\kern-0.50003pth\,$ data and requires no fitting parameters. Furthermore, we combined our low $O\kern-0.50003pth\,$ data with that from previous studies and validated equation (30) from $W\kern-1.49994pte\,=10^{-4}$ to $W\kern-1.49994pte\,=10^{2}$, above which the data begins following a trend closer to $\beta-1\sim W\kern-1.49994pte\,^{1/4}$. As $O\kern-0.50003pth\,$ increases, (30) is less accurate; however, the KM model continues to provide accurate predictions. The KM model also captures the maximum vertical deformation $1-\alpha$ across all $O\kern-0.50003pth\,$, and our energy argument (equation (31)) provides a simple and accurate prediction when $O\kern-0.50003pth\,$ and $B\kern-0.50003pto\,$ effects are negligible. In addition to relating vertical and horizontal deformation, our energy argument allows us to relate maximum deformation and spreading. For example, equation (33) predicts a linear relationship between the horizontal deformation and maximum spread area $\beta_{c}^{2}$, which holds for all $O\kern-0.50003pth\,$ and $B\kern-0.50003pto\,$ tested. This relationship provides a useful tool for determining the contact area of a bouncing drop from its deformation, which is critical for applications such as heat transfer (Shiri & Bird, 2017) or predicting the net charge acquired by a rebounding water drop (Díaz et al., 2022).
@@ 6 Conclusions
We studied a drop rebounding from a non-wetting, rigid substrate at low $W\kern-1.49994pte\,$ using experiments, direct numerical simulation, and reduced-order modeling. We revealed unique rebound behavior at low $W\kern-1.49994pte\,$ that coincides with a change in the retraction dynamics, and showed the critical role of $O\kern-0.50003pth\,$ and $B\kern-0.50003pto\,$ in suppressing rebound. The maximum drop deformation and spreading were connected through energy arguments, and closely followed the predictions of the KM model irrespective of $O\kern-0.50003pth\,$. These findings reveal new physical insight into drop rebound and provide practical tools for predicting drop deformation on non-wetting substrates.
@@
The interplay between our theoretical approaches, the KM model and the dedicated direct numerical simulation setup, enabled us to cross-validate these tools and establish the range of validity of the reduced-order model with a significant level of confidence, with the experimental dataset acting as foundation where comparisons were possible. Thus we were suitably equipped to use the more efficient methodology in the relevant section of the parameter space and relevant distinguished limits thereof, while also being able to rely on DNS for larger values of $W\kern-1.49994pte\,$ where the model assumptions no longer hold. Furthermore, our work also presents the first application of the kinematic match method to the case of a deformable impactor rebounding off of a rigid substrate. The quality of the resulting predictions highlights the versatility of the kinematic match method and naturally invites further applications for it; e.g., its use in the case of a collision of two deformable objects, and in combination with moving meshes and variational methods.
@@
When contact line formation is avoided the coefficient of restitution $\varepsilon$ is $W\kern-1.49994pte\,$-independent at low $W\kern-1.49994pte\,$ in the inertio-capillary limit. Finite $B\kern-0.50003pto\,$ effects suppress bouncing while increasing $O\kern-0.50003pth\,$ decreases $\varepsilon$ and increases the critical Weber number $W\kern-1.49994pte\,_{c}$ where rebound ceases. In contrast, the contact time $t_{c}$ increases as $W\kern-1.49994pte\,$ decreases at a rate that increases with $O\kern-0.50003pth\,$. The unique trends in $\varepsilon$ and $t_{c}$ at low $W\kern-1.49994pte\,$ coincide with a transition in the spreading and retraction dynamics. In particular, the nearly elastic rebound observed in the inertio-capillary limit correspond to a symmetric spreading and retraction process, while the decreasing $\varepsilon$ at higher $W\kern-1.49994pte\,$ is accompanied by an increasing asymmetry in the spreading and retraction process. Thus, the transition from low- to high-$W\kern-1.49994pte\,$ rebound trends occur once the drop can no longer retract as quickly as it spreads, which depends upon $W\kern-1.49994pte\,$ and $O\kern-0.50003pth\,$ (Bartolo et al., 2005). Using energy arguments we derive simple equations that accurately predict the maximum vertical and horizontal deformation, and the maximum spreading for low $O\kern-0.50003pth\,$ impacts with no fitting parameters. The prediction for the maximum horizontal deformation (30) reveals that the horizontal deformation of an impacting drops scales as $W\kern-1.49994pte\,^{1/2}$, which when compared with experimental data holds up to $W\kern-1.49994pte\,\approx 10^{2}$. Additionally, this simple model accurately relates $\beta-1$ to the maximum vertical deformation $1-\alpha$ via equation (32) and maximum contact area $\beta_{c}^{2}$ via equation (33). For all $O\kern-0.50003pth\,$, the kinematic match model provides excellent predictions for the maximum spreading and deformation for low $W\kern-1.49994pte\,$ rebound where direct numerical simulation struggles due to numerical stiffness and parasitic capillary currents.
@@
Although we focus here on the normal impact of a drop on a rigid surface, the ubiquity of drop impacts make our findings relevant to many situations. For example, low-$W\kern-1.49994pte\,$ rebound governs the final stage of drop dispersal in many applications involving low-$B\kern-0.50003pto\,$ drops, such as agricultural sprays (Kooij et al., 2018; Makhnenko et al., 2021; Chen & Ashgriz, 2022) and cough ejecta (Dbouk & Drikakis, 2020; Bourouiba, 2021), where understanding the spread of the drops contents is essential. Future work is needed to extend our results to more practical and natural settings, which often involve non-Newtonian drops and complaint substrates that have a pronounced effect on impact at high $W\kern-1.49994pte\,$ (Bartolo et al., 2007; Basso & Bostwick, 2020) but remain unexplored for drop rebound at low Weber number. Furthermore, our focus herein has been strictly on the liquid dynamics whereas the coupled air layer dynamics are also a source of rich physics, and have become of increasing interest to modelers Phillips & Milewski (2024); Sprittles (2024).
@@
The authors gratefully acknowledge the financial support of the National Science Foundation (NSF CBET-2123371) and the Engineering and Physical Sciences Research Council (EPSRC EP/W016036/1). C.T.G. gratefully acknowledges the Hope Street Fellowship for additional financial support.

### v2

@@ 4.3 Maximum spreading and deformation
We next examine how drop deformation depends on the governing dimensionless parameters and relate deformation and spreading. To do this, we define the maximum equatorial deformation $\beta-1$, where $\beta=R_{m}/R$ is the nondimensional maximum equatorial radius and $R_{m}=\textrm{max}(r_{m}(t))$. Similarly, we define maximum vertical deformation $1-\alpha$, where $\alpha=H_{min}/2R$ and $H_{min}=\textrm{min}(H(t))$. Lastly, we define the maximum spreading radius as $\beta_{c}=R_{c}/R$, where $R_{c}=\textrm{max}(r_{c}(t))$.
@@ 4.3.1 Energy argument
with the surface energy in the $l=2$ mode being $E_{2}=\frac{8}{5}\pi\sigma\mathscr{A}_{2}^{\prime 2}$ (Moláček & Bush, 2012). Neglecting gravity ($B\kern-0.50003pto\,^{2}\ll W\kern-1.49994pte\,$) and viscous dissipation ($O\kern-0.50003pth\,\ll 1$), and assuming the initial kinetic energy of the droplet ($E_{V}=\frac{2}{3}\pi\rho R^{3}V^{2}$) is fully converted into surface energy of the $l=2$ mode (i.e. $E_{V}=E_{2}$), one finds
@@
using the relations $\mathscr{A}_{2}^{\prime}=2R\left(1-\beta\right)$ and $\mathscr{A}_{2}^{\prime}=R\left(\alpha-1\right)$, respectively, and assuming $\mathscr{A}_{2}^{\prime}\leq 0$ (i.e. $R_{m}\geq R$, $H_{min}\leq 2R$) in the deformed state. The result for $\beta-1$ (equation 30) is identical to that derived by Richard & Quéré (2000) using similar arguments. This derivation also suggests a geometric relationship between $\beta$ and $\alpha$:
@@
Note that equations (30)–(34) have limited predictive power compared with the KM model, which captures viscous effects and the complete evolution of the drop shape. Nevertheless, the favorable comparison between the simple energy argument and KM model elucidates the dominant mode of deformation and motivates the parametric scalings.
@@ 4.3.2 Comparisons
Figure 8: ($a$) Maximum equatorial deformation $\beta-1$ ($\beta=R_{m}/R$) and ($b$) maximum vertical deformation $1-\alpha$ ($\alpha=H_{min}/2R$) plotted against Weber number $W\kern-1.38742pte\,$. Circles represent experiment and stars denote DNS for $B\kern-0.46252pto\,=0.02$. Solid curves show the KM model with $B\kern-0.46252pto\,=0.02$, and the inertio-capillary limit ($O\kern-0.46252pth\,=0$; $B\kern-0.46252pto\,=0$) is represented by black dashed curves. The red dot-dashed curves are predictions from our simplified model (§ 4.3.1). The inset in ($a$) shows our low-viscosity data $O\kern-0.46252pth\,<0.05$ (colored circles) with prior results (diamond (Okumura et al., 2003), pentagon (Clanet et al., 2004), square (Biance et al., 2006), triangle (Laan et al., 2014)). The red dot-dashed curve is equation (30), and the black solid line with slope 1/4 is used to guide the eye. The inset in ($c$) plots $1-\alpha$ against $\beta-1$ for all experiments and the red dot-dashed curve is equation (32). ($c$) Dimensionless spreading area $\beta_{c}^{2}$ against $\beta-1$ for all experiments and DNS results, with equation (33) shown as a red dot-dashed curve. The color bar maps $O\kern-0.46252pth\,$ to color on a logarithmic scale.
@@ 5 Discussion
The coefficient of restitution $\varepsilon$ and contact time $t_{c}$ have different $W\kern-1.49994pte\,$-dependencies at low and high $W\kern-1.49994pte\,$. When $W\kern-1.49994pte\,\gtrapprox 0.5$, the coefficient of restitution $\varepsilon$ decreased as $W\kern-1.49994pte\,$ increased, as observed in prior studies (Biance et al., 2006; Jha et al., 2020). When $W\kern-1.49994pte\,\lessapprox 0.5$, $\varepsilon$ reached a maximum that depends upon $O\kern-0.50003pth\,$ and $B\kern-0.50003pto\,$, and becomes $W\kern-1.49994pte\,$-independent in the inertio-capillary limit. A gravity-induced roll-off in $\varepsilon$ leading to $\varepsilon\rightarrow 0$ is the natural route to rebound suppression when $B\kern-0.50003pto\,$ is finite, with $O\kern-0.50003pth\,$ increasing the critical Weber number $W\kern-1.49994pte\,_{c}$ below which rebound stops. These results differ from the peak and subsequent roll-off in $\varepsilon$ as $W\kern-1.49994pte\,$ decreases reported for $W\kern-1.49994pte\,<1$ rebounds on superhydrophobic substrates (Richard & Quéré, 2000; Wang et al., 2022; Thenarianto et al., 2023). Although these prior experiments were conducted in low $O\kern-0.50003pth\,$ conditions, and in some cases used low $B\kern-0.50003pto\,$ drops, the increased effect of contact line friction at low $W\kern-1.49994pte\,$ was shown to be the primary mechanism for rebound suppression. Consequently, our results show that submillimeter drops, like those from coughing (Dbouk & Drikakis, 2020) or agricultural sprays (Kooij et al., 2018; Makhnenko et al., 2021; Chen & Ashgriz, 2022), may continue bouncing longer than previously thought when the substrate is non-wetting—an important insight into the spread of their contents. The contact time between the drop and substrate is $W\kern-1.49994pte\,$-dependent at low $W\kern-1.49994pte\,$ (Foote, 1975; Okumura et al., 2003), and our results show that it is also $O\kern-0.50003pth\,$-dependent in this regime. Increasing $O\kern-0.50003pth\,$, causes $t_{c}$ to increase at a higher rate as $W\kern-1.49994pte\,$ decreases such that viscosity is most influential on $t_{c}$ in the low-$W\kern-1.49994pte\,$ regime. Since $t_{c}$ is a critical parameter governing drop–substrate interaction in various applications Shiri & Bird (2017); D´ıaz et al. (2022), this finding has immediate practical implications, and motivates the study of low $W\kern-1.49994pte\,$ rebound using drops with shear-dependent viscosity in future work.
@@
In section 4.2, we compared the drop shape during spreading and retraction for three values of $W\kern-1.49994pte\,$ spanning the transition from low $W\kern-1.49994pte\,$ to high $W\kern-1.49994pte\,$, and revealed a nearly symmetric spreading and retraction process for low $O\kern-0.50003pth\,$, low $W\kern-1.49994pte\,$ rebound. As $W\kern-1.49994pte\,$ increased, the spread time decreased relative to the retraction time, increasing asymmetry between spreading and retraction and aligning with the changes in rebound behavior. This change corresponds to a loss of energy during rebound when $W\kern-1.49994pte\,>1$ and was investigated through a careful energy budget analysis by Gilet & Bush (2012). Their results showed that when $W\kern-1.49994pte\,<1$, the inital translational energy was efficiently converted to surface energy during the spreading phase. When $W\kern-1.49994pte\,>1$, this conversion was inefficient, with an appreciable loss of energy during spreading. This suggests that the time symmetry is broken in the spreading phase at high $W\kern-1.49994pte\,$. Additionally, Bartolo et al. (2005) identified a limit on the retraction rate of a drop that is substrate-independent for low $O\kern-0.50003pth\,$. Although this limit does not explain the onset of asymmetry, which requires considering the energy lost during spreading, it helps rationalize how the spreading-to-retraction time ratio evolves as $W\kern-1.49994pte\,$ increases when $W\kern-1.49994pte\,>1$. In our experiments, this ratio follows the scaling $t_{s}/t_{r}\sim W\kern-1.49994pte\,^{-0.26\pm 0.01}$ in the low $O\kern-0.50003pth\,$ limit. We can rationalize this observation by considering the inertio-capillary limit, where the time needed for a drop to spread $t_{s}\sim R_{c}/V$, where $R_{c}\sim R\,W\kern-1.49994pte\,^{-1/4}$ from our energy argument (34), and the maximum retraction rate $t_{r}^{-1}\sim t_{\sigma}^{-1}$ (Bartolo et al., 2005). This analysis gives $t_{s}/t_{r}\sim W\kern-1.49994pte\,^{-1/4}$, in good agreement with our results. We therefore interpret the growing asymmetry in rebound dynamics at $W\kern-1.49994pte\,>1$ as the result of two effects: a loss of energy during spreading that breaks time symmetry early in the rebound process (Gilet & Bush, 2012), and a limited retraction rate that further skews the rebound timescale (Bartolo et al., 2005). Recent studies have also noted the near-symmetry between spreading and retraction times low $W\kern-1.49994pte\,$ (Lin et al., 2024). Our results confirm and extend this observation, by demonstrating that time symmetry holds in the inertio-capillary regime and is progressively lost as either $W\kern-1.49994pte\,$ or $O\kern-0.50003pth\,$ increases.
@@
Lastly, we examined the drop deformation and spreading in section 4.3, and showed that the KM model captures drop deformation well across a wide range of $O\kern-0.50003pth\,$, and agrees with a simple energy model in the low $O\kern-0.50003pth\,$ limit, indicating that deformation is primarily in the $l=2$ mode. The literature on the maximum horizontal deformation is vast and two competing scaling arguments have gained support: one based on momentum conservation $\beta\sim W\kern-1.49994pte\,^{1/4}$ (Clanet et al., 2004; Bartolo et al., 2005; Biance et al., 2006; Tsai et al., 2011; Garc´ıa-Geijo et al., 2020) and the other arising from energy conservation $\beta\sim We^{1/2}$ (Chandra & Avedisian, 1991; Bennett & Poulikakos, 1993; Okumura et al., 2003; Eggers et al., 2010; Villermaux & Bossa, 2011; Laan et al., 2014; Lee et al., 2016). Despite extensive efforts, the correct scaling remains debated, in part due to most studies using the maximum equatorial radius $\beta=R_{m}/R$ to describe horizontal deformation. However, as discussed in (Josserand & Thoroddsen, 2016), $\beta$ ignores the initial drop radius and is inadequate to discern scaling arguments at low $W\kern-1.49994pte\,$. In this work we used the maximum equatorial deformation $\beta-1$, and energy arguments predict $\beta-1=\left(5/48\right)^{1/2}\,We^{1/2}$, as in Richard & Quéré (2000). This prediction is in excellent agreement with our low $O\kern-0.50003pth\,$ data and requires no fitting parameters. Furthermore, we combined our low $O\kern-0.50003pth\,$ data with that from previous studies and validated equation (30) from $W\kern-1.49994pte\,=10^{-4}$ to $W\kern-1.49994pte\,=10^{2}$, above which the data begins following a trend closer to $\beta-1\sim W\kern-1.49994pte\,^{1/4}$. As $O\kern-0.50003pth\,$ increases, (30) is less accurate; however, the KM model continues to provide accurate predictions. The KM model also captures the maximum vertical deformation $1-\alpha$ across all $O\kern-0.50003pth\,$, and our energy argument (equation (31)) provides a simple and accurate prediction when $O\kern-0.50003pth\,$ and $B\kern-0.50003pto\,$ effects are negligible. In addition to relating vertical and horizontal deformation, our energy argument allows us to relate maximum deformation and spreading. For example, equation (33) predicts a linear relationship between the horizontal deformation and maximum spread area $\beta_{c}^{2}$, which holds for all $O\kern-0.50003pth\,$ and $B\kern-0.50003pto\,$ tested. This relationship provides a useful tool for determining the contact area of a bouncing drop from its deformation, which is critical for applications such as heat transfer (Shiri & Bird, 2017) or predicting the net charge acquired by a rebounding water drop (D´ıaz et al., 2022).
@@ 6 Conclusions
We studied a drop rebounding from a non-wetting, rigid substrate at low $W\kern-1.49994pte\,$ using experiments, direct numerical simulation, and reduced-order modeling. We revealed the rebound behavior of drops at low $W\kern-1.49994pte\,$ for a range of $Oh$, showed how increasing $Oh$ or $We$ breaks the symmetry between the spreading and retraction phases, and elucidated the critical role of $O\kern-0.50003pth\,$ and $B\kern-0.50003pto\,$ in suppressing rebound. The maximum drop deformation and spreading were connected through energy arguments, and closely followed the predictions of the KM model irrespective of $O\kern-0.50003pth\,$. These findings reveal new physical insight into drop rebound and provide practical tools for predicting drop deformation on non-wetting substrates.
@@
The interplay between our theoretical approaches, the KM model and the dedicated direct numerical simulation setup, enabled us to cross-validate these tools and establish the range of validity of the reduced-order model with a significant level of confidence, with the experimental dataset acting as foundation where comparisons were possible. Thus, we were suitably equipped to use the more efficient methodology in the relevant section of the parameter space and relevant distinguished limits thereof, while also being able to rely on DNS for larger values of $W\kern-1.49994pte\,$ where the model assumptions no longer hold. In general, DNS is preferable for $W\kern-1.49994pte\,>1$, while the reduced-order model offers high-accuracy predictions with substantially lower computational cost for $W\kern-1.49994pte\,<1$, but continues to yield reasonable estimates up to $W\kern-1.49994pte\,\approx 10$. Furthermore, our work also presents the first application of the kinematic match method to the case of a deformable impactor rebounding off a rigid substrate. The quality of the resulting predictions highlights its versatility and naturally invites further applications for it; e.g., its use in the case of a collision of two deformable objects, and in combination with moving meshes and variational methods.
@@
When contact line formation is avoided, the coefficient of restitution $\varepsilon$ is $W\kern-1.49994pte\,$-independent at low $W\kern-1.49994pte\,$ in the inertio-capillary limit. Finite $B\kern-0.50003pto\,$ effects suppress bouncing while increasing $O\kern-0.50003pth\,$ decreases $\varepsilon$ and increases the critical Weber number $W\kern-1.49994pte\,_{c}$ where rebound ceases. In contrast, the contact time $t_{c}$ increases as $W\kern-1.49994pte\,$ decreases at a rate that increases with $O\kern-0.50003pth\,$. The shift in how $\varepsilon$ and $t_{c}$ change with $W\kern-1.49994pte\,$ in the low $We$ and high $We$ regimes coincides with altered spreading and retraction dynamics. In particular, the nearly elastic rebound observed in the inertio-capillary limit have a nearly-symmetric spreading and retraction phase, but increasing $We$ above unity and increasing $Oh$ decreases $\varepsilon$ and increases asymmetry in the spreading and retraction process. Using energy arguments, we derive simple equations that accurately predict the maximum vertical and horizontal deformation, and the maximum spreading for low $O\kern-0.50003pth\,$ impacts with no fitting parameters. The prediction for the maximum horizontal deformation (30) reveals that the horizontal deformation of an impacting drops scales as $W\kern-1.49994pte\,^{1/2}$, which when compared with experimental data holds up to $W\kern-1.49994pte\,\approx 10^{2}$. Additionally, this simple model accurately relates $\beta-1$ to the maximum vertical deformation $1-\alpha$ via equation (32) and maximum contact area $\beta_{c}^{2}$ via equation (33). For all $O\kern-0.50003pth\,$, the kinematic match model provides excellent predictions for the maximum spreading and deformation for low $W\kern-1.49994pte\,$ rebound where direct numerical simulation struggles due to numerical stiffness and parasitic capillary currents.
@@
Although we focus here on the normal impact of a drop on a rigid surface, the ubiquity of drop impacts make our findings relevant to many situations. For example, low-$W\kern-1.49994pte\,$ rebound governs the final stage of drop dispersal in many applications involving low-$B\kern-0.50003pto\,$ drops, such as agricultural sprays (Kooij et al., 2018; Makhnenko et al., 2021; Chen & Ashgriz, 2022) and cough ejecta (Dbouk & Drikakis, 2020; Bourouiba, 2021), where understanding the spread of the drops contents is essential. In both examples, the drops typically impact non-wetting surfaces, as plant leaves are often naturally superhydrophobic (Barthlott & Neinhuis, 1997), while medical surfaces are frequently rendered superhydrophobic to reduce microbial adhesion and biofluid contamination (Falde et al., 2016). Future work is needed to extend our results to more complex practical and natural settings, which often involve non-Newtonian drops and complaint substrates that have a pronounced effect on drop impact at high $W\kern-1.49994pte\,$ (Bartolo et al., 2007; Basso & Bostwick, 2020) but remain unexplored for drop rebound at low Weber number. Furthermore, our focus herein has been strictly on the liquid dynamics whereas the coupled air layer dynamics are also a source of rich physics, and have become of increasing interest to modelers Phillips & Milewski (2024); Sprittles (2024).
@@
The authors gratefully acknowledge the financial support of the National Science Foundation (NSF CBET-2123371) and the Engineering and Physical Sciences Research Council (EPSRC EP/W016036/1). C.T.G. gratefully acknowledges the Hope Street Fellowship for financial support.